A Complete Bibliography of ACM Transactions on Office Information Systems

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 November 2011 Version 1.06

Title word cross-reference

1 [BLT84].

Access [FC84, RS88]. Accessibility [Cul84]. Accounting [BGMW83]. Acquisition [TFGN87]. Act [ALL88]. Action [Suc83]. Activities [HS84]. Adaptation [BBG85]. Addressing [Hau86, Tsi84]. Admissibility [KS85]. Adoption [Ehr87]. Agent [OD83]. AI [SJ85]. Analysis [CEG^{+}87, FC87, HKF84]. Analytic [Hir86]. Analytical [FC84]. Analyzing [GS86, Pan84]. Applications [BCG^{+}87, BFG85, Kel84, TA84]. Appraisal [HS84]. Approach [ALL88, GT83, GB83, HB85, TW86]. Architecture [ABB^{+}84, HM85, PJN85].

Area [HWY85]. ASK [TT85]. Assessment [KDK85]. Assistants [KK87]. Augmentation [CEG^{+}87]. Augmenting [RM87]. Authentication [IL83]. Automated [SB83]. Automation [PS84].

BAROQUE [Mot86a]. Based [ALL88, BLT84, CL87, KK87, Nec88, Pol88, TW86]. Behavioral [GS87a]. Benefit [Sas87]. Between [Nec88]. Bidirectional [BBG85]. book [Suc88]. Browser [Mot86a]. Bureaucracies [Lee88]. Business [CG87]. Butler [DH85].

Calendars [KDK85]. Certain [Pan85]. Challenges [GB83]. Channel [TK84]. Clearinghouse [OD83]. Co [BJ86]. Co-oP [BJ86]. Cognitive [HB85]. Command [AS83, HKF84]. Command-Language [AS83]. Communicating [Tri88]. Communication [Ehr87, TA84, TK84]. Communications [BJ86, Nic85, SB83]. Complete [MT86]. Computer [EB88, FGHW88, Gre88b, KS85, MGL^{+}87, MT86]. Computer-Generated [KS85]. Computer-Supported [Gre88b, MGL^{+}87]. Computing [Gas86]. Concept [DS87]. Conference [HZ86b]. Conferences [Gre88b]. Contexts [DS87]. Control [BGMW83, PS84]. Cooperating [KK87]. Cooperative [Gre88b, LMY88]. Coordination [Hol88, MGL^{+}87]. COSIE [TA84]. Cost [Sas87]. Cost-Benefit [Sas87]. Court [KS85]. Creating [BG87, GB83]. Critical [HS84]. Current [KDK85]. Cursor [GLB85]. Customization [Dam85].

Data [BCG^{+}87, GT83, GS87b, Nic85]. Database [Dam85, DZ86, DW86, FBC^{+}87, HZ87, KM85, SSL^{+}83, WAB^{+}87]. Databases [Mot86a, Mot88]. Datalog [HG85]. Decentralized [OD83]. Decision [BJ86, TK84]. Demand [SB83]. Deontic [Lee88]. Department [CG87]. Description [FC87]. Design [BP84, BJ86, FGHW88, HKF84, HB85, Kel84, KM85, Mal83, PFK83, Suc83]. Designers [Nec88]. Designing [Suc88]. Desks [Mal83]. Details [AS83]. Dimensions [Cul84]. Diplans [Hol88]. Direct [HK86]. Discussion [CB88]. Dissemination [TFGN87]. Distributed [LM84, Nut83, OD83, TA84, WL86]. DITROFF [BBG85]. DITROFF/FFORTID [BBG85]. Diverse [SHC83]. Diversity [Mac88]. Do [Mal83]. Document [BRG88, CTH^{+}86, PJN85, SSL^{+}83, TSW^{+}85]. Documents [FC84]. Domain [MF85]. Dozen [TT85]. Driven [HB85]. Due [GS86]. During [GLB85].

Early [SBF^{+}87]. Editing [GLB85]. Editor [Ell84, Lim83, Win88]. Editorial [All85b, All85a, Bal85, Gre88a, GE87, Loc87]. Editors [HZ86a]. Effective [TK84]. Electronic [KDK85, Mac88, RS88, SB83, TSW^{+}85, WB85]. Electronically [All85a]. Encouraging [Ehr87]. Encyclopedia [WB85]. Ends [Dam85]. Ensuring [KS85]. Environment [OD83, Tri88]. Environments [BLT84]. Etiquette [Bro83]. Evaluation [BGMW83, FC84]. Evolution [CG87]. Example [WAB^{+}87]. Experiences [SBF^{+}87]. Experiment [EB88]. Experimental [BG87, JD86, Nut83, PFKR88]. Experiments [GS87a]. Exploratory [CB88]. Extraction [CTH^{+}86].

Factors [GB83]. Federal [Pan85].

Federated [HM85]. FFORTID [BBG85]. Field [EB88]. File [FC87]. Files [FC84, TC83]. Filing [FC87, GB83]. Filtering [Pol88]. Firm [CG87]. FORMANAGER [YHSL84]. Formation [CTH^{+}86]. Formatting [BBG85, PJN85]. Forms [YHSL84]. Friendly [Kel84]. Front [Dam85].

Generated [KS85]. Generator [HK86]. gIBIS [CB88]. Goals [Nec88]. Government [Pan85]. Graphical [DW86]. Group [BJ86, EB88, GS87b]. Guest [Win88]. Guided [Tri88].

Half [TT85]. Handling [TW86].

Handmarkings [GS87a]. Human [GB83]. Hypertext [CB88, DS87, Tri88].

Images [TSW^{+}85]. Impact [PS84]. Imperative [BFG85]. Implementation [Hol88]. Implementing [CS88, Cul84]. Implications [Cul84, HB85, Mal83]. Individual [Pan84]. Information [CTH^{+}86, Cul84, GT83, HM85, HZ86b, HHK86, Kel84, KM85, LM84, Mal83, ML84, RM87]. Inquiry [Mac88]. Insurance [CG87]. Integrated [Nic85, WAB^{+}87]. Integrating [PSM87, SHC83]. Integration [Gas86]. Intensive [Pan85]. Interaction [FGHW88]. Interactive [BFG85]. Interface [CG87, DZ86, HKF84, Mot88, Nec88]. Interfaces [JD86, SBF^{+}87]. Internal [BGMW83]. Internet [Rad84]. Internetwork [Rad84]. Internetworks [IL83]. Intraoffice [Pah88]. Introduction [Bal85, Ell84, Gre87, Gre88a, GE87, HZ86a, Lim83, Loc87, Win88]. Iris [FBC^{+}87]. Issue [Bal85, Ell84, Gre88a, GE87, HZ86a, Loc87, Win88]. Issues [BCG^{+}87]. Iterative [Kel84].

Keystrokes [AS83]. KNOs [TFGN87]. Knowledge [BLT84, CL87, KK87, Nec88, SHC83, TFGN87]. Knowledge-Based [BLT84, CL87, KK87, Nec88].

Language [AS83, BLT84, Dam85, Eps85, Hol88, Kel84, SJ85]. Languages [SJ85]. Large [CG87]. Laurel [Bro83]. LDC [BLT84]. LDC-1 [BLT84]. Learning [MLC83]. Lens [LMY88]. Linguistic [MF85]. Links [Rad84]. Local [HWY85]. Locating [OD83]. Location [JD86]. Logical [MT86, ML84].

Machines [DH85]. Mail [Mac88, MT86, PFKR88]. Making [TK84]. Management [ABB^{+}84, CL87, DZ86, FBC^{+}87, HM85, LM84, YHSL84]. Manager [WAB^{+}87]. Managerial [TK84]. Managing [Rad84]. Manipulation [HK86, TFGN87]. Manuscripts [All85a]. MCC [SJ85]. Mechanism [Mot86b]. Medical [MF85]. Meetings [CEG^{+}87]. Memory [HZ87]. Message [Bro83, CL87, Pol88, TC83, Tsi84]. Messages [MGL^{+}87]. Messaging [RS88]. Metaphor [JD86]. Method [FC84]. Methodology [Kel84, KM85, Sas87]. Methods [FC87]. MINOS [CTH^{+}86]. Model [BGMW83, BCG^{+}87, CTH^{+}86, Hau86, HHK86]. Modeling [ALL88, GT83, Nut83]. Models [Suc83]. Mores [Bro83]. Movement [GLB85]. Multimedia [BRG88, CTH^{+}86, PFKR88]. Multiuser [SBF^{+}87]. Multivariate [HKF84].

Name [JD86]. Named [OD83]. Naming [Hau86]. Natural [BLT84, Dam85, Eps85, Kel84, SJ85]. Navy [MF85]. Needs [KDK85, Pan84]. Network [TW86]. Network-Based [TW86]. Networks [HWY85, Pah88]. Nick [CEG^{+}87].

Object [ABB^{+}84, BCG^{+}87, FBC^{+}87, HZ87, LM84, PSM87, LMY88]. Object-Oriented [BCG^{+}87, FBC^{+}87, HZ87]. Objects [OD83, TFGN87]. Office [ALL88, BLT84, BFG85, BP84, CL84, Cul84, Ehr87, FC87, GT83, GB83, HB85, HZ86b, HS84, Hir86, HHK86, HK86, IL83, KK87, Kel84, KDK85, Mal83, ML84, PS84, Pan85, Sas87, Suc83, TA84, WAB^{+}87, WL86, YHSL84, WAB^{+}87]. Office-by-Example [WAB^{+}87]. Office-Intensive [Pan85]. Offices [Hew86, Pan84, SB83]. OIS [ABB^{+}84, BGMW83]. Online [Cul84]. oP [BJ86]. Open [Hew86]. Optimization [HWY85]. Organization [PS84]. Organizational [Bar83, CG87, FGHW88]. Organizations [WL86]. Organize [Mal83]. Oriented [BCG^{+}87, FBC^{+}87, HZ87]. Other [PSM87, SJ85]. Outcomes [RS88].

Papers [Gre88b]. Partitioning [DS87]. Patterns [Nic85]. People [BG87, Mal83]. Performance [FC84, FC87]. Permits [Mot88]. Perspective [Hir86]. Planning [PS84]. Policy [CB88]. Portability [HG85]. Practical [Suc83]. PRE [Eps85]. Preliminary [Mac88]. Presentation [CTH^{+}86]. Presuppositions [Mot86b]. Principal [GB83]. Problem [Bar83, WL86].

Problems [Dam85, MLC83]. Procedure [Suc83]. Process [DH85, GS86]. Processing [All85a, BRG88, Eps85, PFK83, SJ85, SSL^{+}83]. Processor [BLT84, BFG85]. Processors [MLC83]. Productivity [Pan85]. Program [SJ85]. Project [MF85, SJ85, CEG^{+}87]. Projecting [SB83]. Projects [PS84]. Promote [Nec88]. Prospects [MLC83]. Prototype [TSW^{+}85, WB85].

Queries [Mot88]. Query [BRG88, HWY85, Mot86b].

Ranking [CS88]. Recognition [SHC83]. Records [KS85, PFK83]. Reference [JD86]. Relational [Mot86a, Mot88, SSL^{+}83]. Requirements [BP84]. Resource [DH85]. Retrieval [TSW^{+}85]. review [Suc88]. Revised [SBF^{+}87]. Routine [Gas86]. Routing [Hau86, ML84]. Routings [MT86]. Rule [Pol88]. Rule-Based [Pol88].

Schemes [Tsi84]. SEAVE [Mot86b]. Section [Gre87]. Sectors [Pan85]. Segmented [HZ87]. Selected [Gre88b]. Selection [TK84]. Semantics [HG85]. Semistructured [MGL^{+}87]. Sentence [BFG85]. Server [PSM87]. Shared [HZ87, Nec88]. Sharing [DH85, GS87b]. Signature [FC84, FC87]. Signatures [CS88]. Simplicity [Eps85]. Social [Hir86]. Social-Analytic [Hir86]. Society [HHK86]. Solutions [Dam85]. Solving [Bar83, WL86]. Some [Dam85]. Sources [SHC83]. Spatial [JD86]. Special [Bal85, Ell84, Gre87, Gre88a, GE87, HZ86a, Loc87, Win88]. Specification [ML84]. Speech [ALL88, GB83]. Speech-Act-Based [ALL88]. Spice [DH85]. Spreadsheet [LMY88]. Spreadsheets [BG87]. Star [PFK83]. Station [Bar83]. Storage [TSW^{+}85]. Strategies [CS88, Ehr87]. String [MF85]. Structures [EB88]. Study [BG87, Hol88]. Studying [HS84]. Subsystem [TA84]. Successful [Ehr87]. Support [BJ86, CL84, EB88, GB83, TA84]. Supported [Gre88b, MGL^{+}87]. Supporting [Bar83, WL86]. Surprisingly [MGL^{+}87]. Syntax [HG85]. System [BRG88, Bro83, BJ86, CL87, CTH^{+}86, CL84, Eps85, FBC^{+}87, GB83, HZ87, IL83, MF85, Nic85, Nut83, Pol88, PFKR88, RS88, SSL^{+}83, Suc83, TSW^{+}85, WAB^{+}87, YHSL84]. Systems [BP84, Cul84, Ehr87, FGHW88, GT83, HZ86b, Hew86, HHK86, HK86, KM85, Lee88, LM84, Mal83, MT86, ML84, Mot86b, RM87, Sas87].

Tabletops [Tri88]. Task

[CL84, HB85, HS84]. Task-Driven [HB85]. Tasks [HB85]. Taxonomies [HS84]. Technology [KDK85]. Terminology [Nec88]. Tests [JD86]. Text [BBG85, CS88, GLB85, PJN85, SHC83, TW86]. TEXTNET [TW86]. Their [Mal83, PS84]. Thesauri [RM87]. Tool [CB88, DW86, HS84, KM85]. Tools [Nec88, Tri88]. Tours [Tri88]. Transient [Rad84]. Transportability [SJ85]. Transportable [BLT84, Eps85, TT85]. Transporting [MF85]. Trends [Pan85].

U.S. [Pan85]. Understanding [Hir86]. UNIX [BBG85, HKF84]. Unstructured [HB85]. Usage [Nic85, RS88]. Use [HKF84, MLC83, Mac88]. Useful [MGL^{+}87]. User [JD86, Kel84, KDK85, Mot86b, Mot88, Suc88]. User-Friendly [Kel84]. Using [CS88].

Vague [Mot88, Mot88]. Verifying [Mot86b]. Version [DZ86]. versus [JD86]. Visual [DZ86]. Voice [BFG85, Nic85].

Ways [TT85]. Whiteboards [DW86]. Wireless [Pah88]. Word [MLC83]. Work [Bar83, EB88, Gas86, GS87b, Gre88b, HB85, LMY88, Suc83]. Workplace [GS86]. Worlds [PSM87]. WYSIWIS [SBF^{+}87].

Xerox [Rad84].

References

Ahlsen:1984:AOM

[ABB^{+}84] M. Ahlsen, A. Bjornerstedt, S. Britts, C. Hulten, and L. Soderlund. An architecture for object management in OIS. ACM Transactions on Office Information Systems, 2(3):173–196, July 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Research Con-

tributions: Selected Papers from the Conference on Office Information Systems.

Allen:1985:EPM

[All85a] R. B. Allen. Editorial: Processing manuscripts electronically. ACM Transactions on Office Information Systems, 3(3):233, July 1985. CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.
org:80.

Allen:1985:E

[All85b] Robert B. Allen. Editorial. ACM Transactions on Office Information Systems, 3(1):1, January 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Auramaki:1988:SAB

[ALL88] Esa Auramaki, Erkki Lehtinen, and Kalle Lyytinen. A speechact-based office modeling approach. ACM Transactions on Office Information Systems, 6 (2):126–152, April 1988. CO-DEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Special Issue on the Language/Action Perspective.

Allen:1983:DCL

[AS83] R. B. Allen and M. W. Scerbo. Details of command-language keystrokes. ACM Transactions on Office Information Systems, 1 (2):159–178, April 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Ballard:1985:EIO

[Bal85] B. W. Ballard. Editorial: Introduction to the special issue. ACM Transactions on Office Information Systems, 3(2):105–106, April 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Barber:1983:SOP

[Bar83] G. Barber. Supporting organizational problem solving with a work station. ACM Transactions on Office Information Systems, 1 (1):45–67, January 1983. CO-DEN ATOSDO. ISSN 0734-2047.

URL http://www.acm.org:80.

Buchman:1985:DFA

[BBG85] C. Buchman, D. M. Berry, and J. Gonczarowski. DITROFF/ FFORTID, an adaptation of the UNIX DITROFF for formatting bidirectional text. ACM Transactions on Office Information Systems, 3(4):380–397, October 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Banerjee:1987:DMI

[BCG^{+}87] Jay Banerjee, Hong-Tai Chou, Jorge F. Garza, Won Kim, Darrell Woelk, Nat Ballou, and Hyoung-Joo Kim. Data model issues for object-oriented applications. ACM Transactions on Office Information Systems, 5 (1):3–26, January 1987. CO-DEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Special Issue on Object-Oriented Systems.

Biermann:1985:ISP

[BFG85] A. W. Biermann, L. Fineman, and K. C. Gilbert. An imperative sentence processor for voice interactive office applications. ACM Transactions on Office Information Systems, 3(4):321–346, October 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Brown:1987:ESP

[BG87] Polly S. Brown and John D.

Gould. An experimental study of people creating spreadsheets. ACM Transactions on Office Information Systems, 5(3):258–272, July 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Bailey:1983:OMI

[BGMW83] A. D. Bailey, J. H. Gerlach, R. P.

McAfee, and A. B. Whinston. An OIS model for internal accounting control evaluation. ACM Transactions on Office Information Systems, 1(1):25–44, January 1983. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Bui:1986:CDC

[BJ86] T. X. Bui and M. Jarke. Communications design for Co-oP: A group decision support system. ACM Transactions on Office Information Systems, 4(2):81–103, April 1986. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Ballard:1984:LTK

[BLT84] B. W. Ballard, J. C. Lusth, and N. L. Tinkham. LDC-1: A transportable, knowledge-based natural language processor for office environments. ACM Transactions on Office Information Systems, 2(1):1–25, January 1984. CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.
                                        org:80.

Bracchi:1984:DRO

[BP84] G. Bracchi and B. Pernici. The design requirements of office systems. ACM Transactions on Office Information Systems, 2(2): 151–170, April 1984. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

Bertino:1988:QPM

[BRG88] Elisa Bertino, Fausto Rabitti, and Simon Gibbs. Query processing in a multimedia document system. ACM Transactions on Office Information Systems, 6(1):

1–41, January 1988. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Brotz:1983:MSM

[Bro83] D. K. Brotz. Message system mores: Etiquette in Laurel. ACM Transactions on Office Information Systems, 1(2):179–192, April 1983. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Conklin:1988:GHT

[CB88] Jeff Conklin and Michael L.

Begeman. gIBIS: A hypertext tool for exploratory policy discussion. ACM Transactions on Office Information Systems, 6(4): 303–331, October 1988. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80. Se-

lected Papers from the Conference on Computer-Supported Cooperative Work (CSCW ’88).

Cook:1987:PNM

[CEG^{+}87] Peter Cook, Clarence Ellis, Mike Graf, Gail Rein, and Tom Smith. Project Nick: Meetings augmentation and analysis. ACM Transactions on Office Information Systems, 5(2):132–146, April 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Special Issue

on Computer-Supported Cooperative Work.

Clement:1987:EOI

[CG87] Andrew Clement and C. C.

Gotlieb. Evolution of an organizational interface: the new business department at a large insurance firm. ACM Transactions on Office Information Systems, 5 (4):328–339, October 1987. CO-DEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.

org:80. Special Section: Selected Papers from CHI+GI ’87.

Croft:1984:TSO

[CL84] W. B. Croft and L. S. Lefkowitz. Task support in an office system. ACM Transactions on Office Information Systems, 2(3):197–212, July 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80. Research Con-

tributions: Selected Papers from the Conference on Office Information Systems.

Chang:1987:KBM

[CL87] Shi-Kuo Chang and L. Leung. A knowledge-based message management system. ACM Transactions on Office Information Systems, 5(3):213–236, July 1987. CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.
                                        org:80.

Croft:1988:IRS

[CS88] W. Bruce Croft and Pasquale Savino. Implementing ranking strategies using text signatures. ACM Transactions on Office Information Systems, 6(1): 42–62, January 1988. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

Christodoulakis:1986:MDP

[CTH^{+}86] S. Christodoulakis, M. Theodoridou, F. Ho, M. Papa, and A. Pathria. Multimedia document presentation, information extraction, and document formation in MINOS: A model and a system. ACM Transactions on Office Information Systems, 4(4): 345–383, October 1986. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Culnan:1984:DAO

[Cul84] M. J. Culnan. The dimensions of accessibility to online information: Implications for implementing office information systems. ACM Transactions on Office Information Systems, 2(2):141–150, April 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Damerau:1985:PSS

[Dam85] F. J. Damerau. Problems and some solutions in customization of natural language database front ends. ACM Transactions on Office Information Systems, 3 (2):165–184, April 1985. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Dannenberg:1985:BPR

[DH85] R. B. Dannenberg and P. G. Hibbard. A Butler process for resource sharing on Spice machines. ACM Transactions on Office Information Systems, 3(3):234–252, July 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Delisle:1987:CPC

[DS87] Norman M. Delisle and Mayer D. Schwartz. Contexts — A partitioning concept for hypertext. ACM Transactions on Office Information Systems, 5(2):168–186, April 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80. Special Issue

on Computer-Supported Cooperative Work.

Donahue:1986:WGD

[DW86] James Donahue and Jennifer Widom. Whiteboards: A graphical database tool. ACM Transactions on Office Information Systems, 4(1):24–41, January 1986. CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.
                                        org:80.

Davison:1986:VID

[DZ86] Jay W. Davison and Stanley B. Zdonik. A visual interface for a database with version management. ACM Transactions on Office Information Systems, 4(3): 226–256, July 1986. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80. Re-

search Contributions: Selected Papers from the Conference on Office Information Systems.

Eveland:1988:WGS

[EB88] J. D. Eveland and T. K. Bikson. Work group structures and computer support: A field experiment. ACM Transactions on Office Information Systems, 6(4): 354–379, October 1988. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80. Se-

lected Papers from the Conference on Computer-Supported Cooperative Work (CSCW ’88).

Ehrlich:1987:SES

[Ehr87] Susan F. Ehrlich. Strategies for encouraging successful adoption of office communication systems. ACM Transactions on Office Information Systems, 5(4): 340–357, October 1987. CO-DEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Special Section: Selected Papers from CHI+GI ’87.

Ellis:1984:EIO

[Ell84] C. Ellis. Editor’s introduction to the special issue. ACM Transactions on Office Information Systems, 2(3):171–172, July 1984. CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Research Contributions: Selected Papers from the Conference on Office Information Systems.

Epstein:1985:TNL

[Eps85] S. S. Epstein. Transportable natural language processing through simplicity — the PRE system. ACM Transactions on Office Information Systems, 3(2):107–120, April 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Fishman:1987:IOO

[FBC^{+}87] D. H. Fishman, D. Beech, H. P.

Cate, E. C. Chow, T. Connors, J. W. Davis, N. Derrett, C. G.

Hoch, W. Kent, P. Lyngbaek, B. Mahbod, M. A. Neimat, T. A. Ryan, and M. C. Shan. Iris: An object-oriented database management system. ACM Transactions on Office Information Systems, 5(1):48–69, January 1987. CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Special Issue on Object-Oriented Systems.

Faloutsos:1984:SFA

[FC84] C. Faloutsos and S. Christodoulakis. Signature files: An access method for documents and its analytical performance evaluation. ACM Transactions on Office Information Systems, 2(4):267–288, October 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Faloutsos:1987:DPA

[FC87] Christos Faloutsos and Stavros Christodoulakis. Description and performance analysis of signature file methods for office filing. ACM Transactions on Office Information Systems, 5(3):237–257, July 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Flores:1988:CSD

[FGHW88] Fernando Flores, Michael Graves, Brad Hartfield, and Terry Winograd. Computer systems and the design of organizational interaction. ACM Transactions on Office Information Systems, 6(2):153–172, April 1988. CO-DEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Special Issue on the Language/Action Perspective.

Gasser:1986:ICR

[Gas86] Les Gasser. The integration of computing and routine work. ACM Transactions on Office Information Systems, 4(3):205–225, July 1986. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Research Con-

tributions: Selected Papers from the Conference on Office Information Systems.

Gould:1983:HFC

[GB83] J. D. Gould and S. J. Boies. Human factors challenges in creating a principal support office system — the speech filing system approach. ACM Transactions on Office Information Systems, 1(4): 273–298, October 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Grief:1987:EIS

[GE87] I. Grief and C. Ellis. Editorial:

Introduction to the special issue. ACM Transactions on Office Information Systems, 5(2):113–114, April 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Special Issue

on Computer-Supported Cooperative Work.

Gould:1985:CMD

[GLB85] J. D. Gould, C. Lewis, and V. Barnes. Cursor movement during text editing. ACM Transactions on Office Information Systems, 3(1):22–34, January 1985.

CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.
                                        org:80.

Greif:1987:ISS

[Gre87] I. Greif. Introduction to the special section. ACM Transactions on Office Information Systems, 5(4):327, October 1987. CO-DEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.

org:80. Special Section: Selected Papers from CHI+GI ’87.

Greif:1988:EIS

[Gre88a] I. Greif. Editorial introduction to the special issue. ACM Transactions on Office Information Systems, 6(4):301–302, October 1988. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80. Selected Pa-

pers from the Conference on Computer-Supported Cooperative Work (CSCW ’88).

Greif:1988:SPC

[Gre88b] Irene Greif, editor. Selected Papers from the Conferences on Computer-Supported Cooperative Work, volume 6(4) of ACM Transactions on Office Information Systems. ACM Press, New York, NY 10036, USA, October 1988. CODEN ATISET. ISSN 1046-8188.

Gerson:1986:ADP

[GS86] Elihu M. Gerson and Susan Leigh Star. Analyzing due process in the workplace. ACM Transactions on Office Information Systems, 4(3):257–270, July 1986.

CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Research Contributions: Selected Papers from the Conference on Office Information Systems.

Gould:1987:BEH

[GS87a] John D. Gould and Josiane Salaun. Behavioral experiments on handmarkings. ACM Transactions on Office Information Systems, 5(4):358–377, October 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://www. acm.org:80. Special Section: Selected Papers from CHI+GI ’87.

Greif:1987:DSG

[GS87b] Irene Greif and Sunil Sarin.

Data sharing in group work. ACM Transactions on Office Information Systems, 5(2):187–211, April 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Special Issue

on Computer-Supported Cooperative Work.

Gibbs:1983:DMA

[GT83] S. Gibbs and D. Tsichritzis. A data modeling approach for office information systems. ACM Transactions on Office Information Systems, 1(4):299–319, October 1983. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Hauzeur:1986:MNA

[Hau86] Bernard M. Hauzeur. A model for naming, addressing, and routing. ACM Transactions on Office Information Systems, 4(4):

293–311, October 1986. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

Harris:1985:DIT

[HB85] S. E. Harris and H. J. Brightman. Design implications of a taskdriven approach to unstructured cognitive tasks in office work. ACM Transactions on Office Information Systems, 3(3):292–306, July 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Hewitt:1986:OOS

[Hew86] Carl Hewitt. Offices are open systems. ACM Transactions on Office Information Systems, 4(3): 271–287, July 1986. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80. Re-

search Contributions: Selected Papers from the Conference on Office Information Systems.

Hafner:1985:PSS

[HG85] C. D. Hafner and K. Godden. Portability of syntax and semantics in datalog. ACM Transactions on Office Information Systems, 3(2):141–164, April 1985. CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.
                                        org:80.

Ho:1986:SMO

[HHK86] C.-S. Ho, Y.-C. Hong, and T.- S. Kuo. A society model for office information systems. ACM Transactions on Office Information Systems, 4(2):104–131, April 1986. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Hirschheim:1986:UOS

[Hir86] R. A. Hirschheim. Understanding the office: A social-analytic perspective. ACM Transactions on Office Information Systems, 4(4): 331–344, October 1986. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Hudson:1986:GDM

[HK86] S. E. Hudson and R. King. A generator of direct manipulation office systems. ACM Transactions on Office Information Systems, 4 (2):132–163, April 1986. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Hanson:1984:IDM

[HKF84] S. J. Hanson, R. E. Kraut, and J. M. Farber. Interface design and multivariate analysis of UNIX command use. ACM Transactions on Office Information Systems, 2(1):42–57, January 1984. CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.
org:80.

Heimbigner:1985:FAI

[HM85] D. Heimbigner and D. McLeod.

A federated architecture for information management. ACM Transactions on Office Information Systems, 3(3):253–278, July 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Holt:1988:DNL

[Hol88] Anatol W. Holt. Diplans: A new language for the study and implementation of coordination. ACM Transactions on Office Information Systems, 6(2):109–125, April 1988. CODEN ATOSDO. ISSN 0734-2047. URL http://www. acm.org:80. Special Issue on the Language/Action Perspective.

Higgins:1984:CAT

[HS84] C. A. Higgins and F. R. Safayeni. A critical appraisal of task taxonomies as a tool for studying office activities. ACM Transactions on Office Information Systems, 2 (4):331–339, October 1984. CO-DEN ATOSDO. ISSN 0734-2047.

URL http://www.acm.org:80.

Hevner:1985:QOL

[HWY85] A. R. Hevner, O. Q. Wu, and S. B. Yao. Query optimization on local area networks. ACM Transactions on Office Information Systems, 3(1):35–62, January 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Hewitt:1986:EIS

[HZ86a] C. Hewitt and S. B. Zdonik. Editors’ introduction: Special issue. ACM Transactions on Office Information Systems, 4(3):183–184, July 1986. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Research Con-

tributions: Selected Papers from the Conference on Office Information Systems.

Hewitt:1986:COI

[HZ86b] Carl Hewitt and Stanley B.

Zdonik, editors. 1986 Conference on Office Information Systems, volume 4(3) of ACM Transactions on Office Information Systems. ACM Press, New York, NY 10036, USA, July 1986. CODEN ATISET. ISSN 1046-8188.

Hornick:1987:SSM

[HZ87] Mark F. Hornick and Stanley B. Zdonik. A shared, segmented memory system for an object-oriented database. ACM Transactions on Office Information Systems, 5(1):70–95, January 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Special Issue

on Object-Oriented Systems.

Israel:1983:AOS

[IL83] J. E. Israel and T. A. Linden. Authentication in office system internetworks. ACM Transactions on Office Information Systems, 1 (3):193–210, July 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Jones:1986:SMU

[JD86] William P. Jones and Susan T.

Dumais. The spatial metaphor for user interfaces: Experimental tests of reference by location versus name. ACM Transactions on Office Information Systems, 4(1): 42–63, January 1986. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Kincaid:1985:ECO

[KDK85] C. M. Kincaid, P. B. Dupont, and A. R. Kaye. Electronic calendars in the office: An assessment of user needs and current technology. ACM Transactions on Office Information Systems, 3(1): 89–102, January 1985. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

Kelley:1984:IDM

[Kel84]  J. F. Kelley.  An iterative

design methodology for userfriendly natural language office information applications. ACM Transactions on Office Information Systems, 2(1):26–41, January 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Kaye:1987:CKB

[KK87] A. Roger Kaye and Gerald M. Karam. Cooperating knowledgebased assistants for the office. ACM Transactions on Office Information Systems, 5(4):297– 326, October 1987. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

King:1985:DDM

[KM85] R. King and D. McLeod. A database design methodology and tool for information systems. ACM Transactions on Office Information Systems, 3(1): 2–21, January 1985. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

King:1985:ECA

[KS85] R. King and C. Stanley. Ensuring the court admissibility of computer-generated records. ACM Transactions on Office Information Systems, 3(4):398– 412, October 1985. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Lee:1988:BDS

[Lee88] Ronald M. Lee. Bureaucracies as deontic systems. ACM Transactions on Office Information Systems, 6(2):87–108, April 1988. CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Special Issue on the Language/Action Perspective.

Limb:1983:EI

[Lim83] J. O. Limb. Editor’s introduction. ACM Transactions on Office Information Systems, 1(1): 1–2, January 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Lyngbaek:1984:OMD

[LM84] P. Lyngbaek and D. McLeod. Object management in distributed information systems. ACM Transactions on Office Information Systems, 2(2):96–122, April 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Lai:1988:OLS

[LMY88] Kum-Yew Lai, Thomas W. Malone, and Keh-Chiang Yu. Object Lens: A ‘spreadsheet’ for cooperative work. ACM Transactions on Office Information Systems, 6(4): 332–353, October 1988. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80. Se-

lected Papers from the Conference on Computer-Supported Cooperative Work (CSCW ’88).

Lochovsky:1987:EIS

[Loc87] F. H. Lochovsky. Editorial: Introduction to the special issue. ACM Transactions on Office Information Systems, 5(1):1–2, January 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://www. acm.org:80. Special Issue on Object-Oriented Systems.

Mackay:1988:DUE

[Mac88] Wendy E. Mackay. Diversity in the use of electronic mail: A preliminary inquiry. ACM Transactions on Office Information Systems, 6(4):380–397, October 1988. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Selected Pa-

pers from the Conference on Computer-Supported Cooperative Work (CSCW ’88).

Malone:1983:HDP

[Mal83] T. W. Malone. How do people organize their desks? Implications for the design of office information systems. ACM Transactions on Office Information Systems, 1 (1):99–112, January 1983. CO-DEN ATOSDO. ISSN 0734-2047.

URL http://www.acm.org:80.

Marsh:1985:TLS

[MF85] E. Marsh and C. Friedman.

Transporting the linguistic string project system from a medical to a navy domain. ACM Transactions on Office Information Systems, 3(2):121–140, April 1985. CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.
org:80.

Malone:1987:SMS

[MGL^{+}87] Thomas W. Malone, Kenneth R.

Grant, Kum-Yew Lai, Ramana Rao, and David Rosenblitt. Semistructured messages are surprisingly useful for computersupported coordination. ACM Transactions on Office Information Systems, 5(2):115–131, April 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://www. acm.org:80. Special Issue on Computer-Supported Cooperative Work.

Mazer:1984:LRS

[ML84] M. S. Mazer and F. H. Lochovsky.

Logical routing specification in office information systems. ACM Transactions on Office Information Systems, 2(4):303–330, October 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Mack:1983:LUW

[MLC83] R. L. Mack, C. H. Lewis, and J. M. Carroll. Learning to use word processors: Problems and prospects. ACM Transactions on Office Information Systems, 1(3): 254–271, July 1983. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

Motro:1986:BBR

[Mot86a] A. Motro. BAROQUE: A browser for relational databases. ACM Transactions on Office Information Systems, 4(2):164–181, April 1986. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

                                        acm.org:80.

Motro:1986:SMV

[Mot86b] Amihai Motro. SEAVE: A mechanism for verifying user presuppositions in query systems. ACM Transactions on Office Information Systems, 4(4):312–330, October 1986. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Motro:1988:VUI

[Mot88] Amihai Motro. VAGUE: A user interface to relational databases that permits vague queries. ACM Transactions on Office Information Systems, 6(3):187–214, July 1988. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

                                        acm.org:80.

Martin:1986:CLR

[MT86] P. Martin and D. Tsichritzis. Complete logical routings in computer mail systems. ACM Transactions on Office Information Systems, 4(1):64–80, January 1986. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Neches:1988:KBT

[Nec88] Robert Neches. Knowledge-based tools to promote shared goals and terminology between interface designers. ACM Transactions on Office Information Systems, 6(3): 215–231, July 1988. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Nicholson:1985:UPI

[Nic85] R. T. Nicholson. Usage patterns in an integrated voice and data communications system. ACM Transactions on Office Information Systems, 3(3):307–314, July 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Nutt:1983:EDM

[Nut83] G. J. Nutt. An experimental distributed modeling system. ACM Transactions on Office Information Systems, 1(2):117–142, April 1983. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Oppen:1983:CDA

[OD83] D. C. Oppen and Y. K. Dalal.

The clearinghouse: A decentralized agent for locating named objects in a distributed environment. ACM Transactions on Office Information Systems, 1(3): 230–253, July 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Pahlavan:1988:WIN

[Pah88] K. Pahlavan. Wireless intraoffice networks. ACM Transactions on Office Information Systems, 6 (3):277–302, July 1988. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Panko:1984:OAN

[Pan84] R. R. Panko. 38 offices: Analyzing needs in individual offices. ACM Transactions on Office Information Systems, 2(3):226–234, July 1984. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Research Con-

tributions: Selected Papers from the Conference on Office Information Systems.

Panko:1985:PTC

[Pan85] R. R. Panko. Productivity trends in certain office-intensive sectors of the U.S. federal government. ACM Transactions on Office Information Systems, 3(4): 370–379, October 1985. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Purvy:1983:DSR

[PFK83] R. Purvy, J. Farrell, and P. Klose. The design of Star’s records processing. ACM Transactions on Office Information Systems, 1(1): 3–24, January 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Postel:1988:EMM

[PFKR88] Jonathan B. Postel, Gregory G. Finn, Alan R. Katz, and Joyce K. Reynolds. An experimental multimedia mail system. ACM Transactions on Office Information Systems, 6(1):63–81, January 1988. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80.

Peels:1985:DAT

[PJN85] A. J. H. M. Peels, N. J. M.

Janssen, and W. Nawijn. Document architecture and text formatting. ACM Transactions on Office Information Systems, 3(4): 347–369, October 1985. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Pollock:1988:RBM

[Pol88] Stephen Pollock. A rule-based message filtering system. ACM Transactions on Office Information Systems, 6(3):232–254, July 1988. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

acm.org:80.

Paddock:1984:OAP

[PS84] C. E. Paddock and R. W.

Scamell. Office automation projects and their impact on organization, planning, and control. ACM Transactions on Office Information Systems, 2(4): 289–302, October 1984. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Purdy:1987:IOS

[PSM87] Alan Purdy, Bruce Schuchardt, and David Maier. Integrating an object server with other worlds. ACM Transactions on Office Information Systems, 5 (1):27–47, January 1987. CO-DEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.

org:80. Special Issue on Object-Oriented Systems.

Radicati:1984:MTI

[Rad84] S. Radicati. Managing transient internetwork links in the Xerox Internet. ACM Transactions on Office Information Systems, 2(3): 213–225, July 1984. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80. Re-

search Contributions: Selected Papers from the Conference on Office Information Systems.

Rada:1987:ATI

[RM87] Roy Rada and Brian K. Martin. Augmenting thesauri for information systems. ACM Transactions on Office Information Systems, 5 (4):378–392, October 1987. CO-DEN ATOSDO. ISSN 0734-2047.

                                        URL http://www.acm.org:80.

Rice:1988:AUO

[RS88] Ronald E. Rice and Donald E. Shook. Access to, usage of, and outcomes from an electronic messaging system. ACM Transactions on Office Information Systems, 6(3):255–276, July 1988. CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.
                                        org:80.

Sassone:1987:CBM

[Sas87] Peter G. Sassone. Cost-benefit methodology for office systems. ACM Transactions on Office Information Systems, 5(3):273–289, July 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Smith:1983:PDE

[SB83] S. A. Smith and R. I. Benjamin.

Projecting demand for electronic communications in automated offices. ACM Transactions on Office Information Systems, 1(3): 211–229, July 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Stefik:1987:WRE

[SBF^{+}87] M. Stefik, D. G. Borrow, G. Foster, S. Lanning, and D. Tatar. WYSIWIS revised: Early experiences with multiuser interfaces. ACM Transactions on Office Information Systems, 5(2):147–167, April 1987. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Special Issue

on Computer-Supported Cooperative Work.

Srihari:1983:IDK

[SHC83] S. N. Srihari, J. J. Hull, and R. Choudhari. Integrating diverse knowledge sources in text recognition. ACM Transactions on Office Information Systems, 1(1): 68–87, January 1983. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Slocum:1985:TOL

[SJ85] J. Slocum and C. F. Justus.

Transportability to other languages: The natural language processing project in the AI program at MCC. ACM Transactions on Office Information Systems, 3(2):204–230, April 1985. CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.
                                        org:80.

Stonebraker:1983:DPR

[SSL^{+}83] M. Stonebraker, H. Stettner, N. Lynn, J. Kalash, and A. Guttman. Document processing in a relational database system. ACM Transactions on Office Information Systems, 1(2): 143–158, April 1983. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80.

Suchman:1983:OPP

[Suc83] L. A. Suchman. Office procedure as practical action: Models of work and system design. ACM Transactions on Office Information Systems, 1(4):320–328, October 1983. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Suchman:1988:DUB

[Suc88] L. Suchman. Designing with the user (book review). ACM Transactions on Office Information Systems, 6(2):173–183, April 1988. CODEN ATOSDO. ISSN 0734-2047. URL http:/

                                        /www.acm.org:80. Special Is-

sue on the Language/Action Perspective. Review of Computers and Democracy: A Scandinavian Challenge, G. Bjerknes, P. Ehn, and M. Kyng, Eds. Gower Press, Brookfield, VT, 1987.

Terry:1984:CCS

[TA84] D. B. Terry and S. Andler. The COSIE communication subsystem: Support for distributed office applications. ACM Transactions on Office Information Systems, 2(2):79–95, April 1984. CO-DEN ATOSDO. ISSN 0734-2047.

URL http://www.acm.org:80.

Tsichritzis:1983:MF

[TC83] D. Tsichritzis and S. Christodoulakis.

Message files. ACM Transactions on Office Information Systems, 1(1):88–98, January 1983. CO-DEN ATOSDO. ISSN 0734-2047.

URL http://www.acm.org:80.

Tsichritzis:1987:KKA

[TFGN87] D. Tsichritzis, E. Fiume, S. Gibbs, and O. Nierstrasz. KNOs: KNowledge acquisition, dissemination, and manipulation objects. ACM Transactions on Office Information Systems, 5(1): 96–112, January 1987. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80. Spe-

cial Issue on Object-Oriented Systems.

Trauth:1984:CSE

[TK84] E. M. Trauth and S. K. Kwan.

Channel selection and effective communication for managerial decision making. ACM Transactions on Office Information Systems, 2(2):123–140, April 1984. CODEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.
org:80.

Trigg:1988:GTT

[Tri88] Randall H. Trigg. Guided tours and tabletops: Tools for communicating in a hypertext environment. ACM Transactions on Office Information Systems, 6(4): 398–414, October 1988. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80. Se-

lected Papers from the Conference on Computer-Supported Cooperative Work (CSCW ’88).

Tsichritzis:1984:MAS

[Tsi84] D. Tsichritzis. Message addressing schemes. ACM Transactions on Office Information Systems, 2 (1):58–77, January 1984. CO-DEN ATOSDO. ISSN 0734-2047.

                                        URL http://www.acm.org:80.

Thoma:1985:PSE

[TSW^{+}85] G. R. Thoma, S. Suthasinekul, F. L. Walker, J. Cookson, and M. Rashidian. A prototype system for the electronic storage and retrieval of document images. ACM Transactions on Office Information Systems, 3(3):279–291, July 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://

                                        www.acm.org:80.

Thompson:1985:ATH

[TT85] B. H. Thompson and F. B. Thompson. ASK is transportable in half a dozen ways. ACM Transactions on Office Information Systems, 3(2):185–203, April 1985. CODEN ATOSDO. ISSN 0734-2047. URL http://www.

                                        acm.org:80.

Trigg:1986:TNB

[TW86] Randall H. Trigg and Mark Weiser. TEXTNET: A networkbased approach to text handling. ACM Transactions on Office Information Systems, 4(1): 1–23, January 1986. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Whang:1987:OEI

[WAB^{+}87] Kyu-Young Whang, Art Ammann, Anthony Bolmarcich, Maria Hanrahan, Guy Hochgesang, Kuan-Tsae Huang, Al Khorasani, Ravi Krishnamurthy, Gary Sockut, Paula Sweeney, Vance Waddle, and Moshe Zloof. Office-by-Example: An integrated office system and database manager. ACM Transactions on Office Information Systems, 5(4): 393–427, October 1987. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Weyer:1985:PEE

[WB85] S. A. Weyer and A. H. Borning. A prototype electronic encyclopedia. ACM Transactions on Office Information Systems, 3(1): 63–88, January 1985. CODEN ATOSDO. ISSN 0734-2047. URL

http://www.acm.org:80.

Winograd:1988:GEI

[Win88] T. Winograd. Guest Editor’s introduction to the special issue. ACM Transactions on Office Information Systems, 6(2):83–86, April 1988. CODEN ATOSDO. ISSN 0734-2047. URL http://

www.acm.org:80. Special Issue

on the Language/Action Perspective.

Woo:1986:SDO

[WL86] Carson C. Woo and Frederick H. Lochovsky. Supporting distributed office problem solving in organizations. ACM Transactions on Office Information Systems, 4(3):185–204, July 1986. CODEN ATOSDO. ISSN 0734-

                                        2047. URL http://www.acm.

org:80. Research Contributions: Selected Papers from the Conference on Office Information Systems.

Yao:1984:FOF

[YHSL84] S. B. Yao, A. R. Hevner, Z. Shi, and D. Luo. FORMANAGER: An office forms management system. ACM Transactions on Office Information Systems, 2(3): 235–262, July 1984. CODEN ATOSDO. ISSN 0734-2047. URL

                                        http://www.acm.org:80. Re-

search Contributions: Selected Papers from the Conference on Office Information Systems.